OpenStudy (anonymous):

valuate 3(2m − 3n) − 2(m + 2n) for m = −9 and n = −7.

OpenStudy (anonymous):

6m - 9n - 2m - 4n 6(-9) - 9(-7) - 2(-9) - 4(-7) = -54 + 63 + 18 + 28 = 55

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

no problem(: